# SIC-POVMs and orders of real quadratic fields

### v1

@@ 3.2. Fields generated by Weyl–Heisenberg SICs
Proof of (1): Let ${\mathbf{v}}=(v_{0},v_{1},\ldots,v_{d-1})^{\top}$ be a fiducial vector for $S$, and let $\Pi={\mathbf{v}}{\mathbf{v}}^{\dagger}$. It has $(1,0)$-entry $(\Pi)_{10}=v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. Now the modulation matrix $Z\in\operatorname{WH}(d)$ gives another SIC vector $Z{\mathbf{v}}=(v_{0},e^{\frac{2\pi i}{d}}v_{1},\cdots,e^{\frac{2\pi i(d-1)}{d}}v_{d-1})$. The corresponding projection $\Pi^{\prime}:=Z{\mathbf{v}}(Z{\mathbf{v}})^{\dagger}=Z{\mathbf{v}}{\mathbf{v}}^{\dagger}Z^{-1}$ has $(\Pi^{\prime})_{1,0}=e^{\frac{2\pi i}{d}}v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. If $v_{1}\overline{v}_{0}\neq 0$, then we may conclude $e^{\frac{2\pi i}{d}}\in F^{\rm proj}\!\left(S\right)$.
We show every Weyl–Heisenberg SIC has a fiducial vector ${\mathbf{v}}$ with both $v_{0}$ and $v_{1}$ nonzero, so the argument applies. We first claim that every fiducial vector ${\mathbf{v}}$ has two consecutive entries $v_{i},v_{i+1}$ nonzero, allowing the wraparound product $v_{d-1}v_{0}$. Suppose not; then, the translation matrix $X\in\operatorname{WH}(d)$ gives a cyclic shift $X{\mathbf{v}}=(v_{1},v_{2},...,v_{d-1},v_{0})$, so $\langle{\mathbf{v}},X{\mathbf{v}}\rangle=\sum_{i=0}^{d-1}v_{i}\overline{v}_{i+1}=0$, contradicting the SIC property $\left|\langle{\mathbf{v}},X{\mathbf{v}}\rangle\right|=\frac{1}{d+1}$. The claim follows. If the given ${\mathbf{v}}$ has $v_{i}v_{i+1}\neq 0$, then ${\mathbf{v}}^{\prime}=X^{d-i}{\mathbf{v}}$ cyclically shifts entries so that $v_{0}^{\prime}v_{1}^{\prime}\neq 0$.
@@
Proof of (2): By (1), $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)(e^{\frac{\pi i}{d}})$ is at most a quadratic extension of $F^{\rm proj}\!\left(S\right)$, since the field index $[\mathbb{Q}(e^{\frac{\pi i}{d}}):\mathbb{Q}(e^{\frac{2\pi i}{d}})]=2$ if $d$ is even and is $1$ if $d$ is odd. ∎
@@
where $X,Z$ are Weyl–Heisenberg group generators, $F$ is the discrete Fourier transform,
$F_{j,k}=\frac{1}{\sqrt{d}}e^{\frac{2\pi ijk}{d}}\,\mbox{ for }\,0\leq j,k\leq d-1.$ (3.10)
and $R$ is a diagonal matrix, with diagonal entries
@@
This is [46, Thm. 14.1]. ∎
@@
Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$. For any $M\in\operatorname{EC}(d)$,
@@
where $MS$ denotes the Weyl–Heisenberg line-SIC having fiducial vector ${\mathbf{w}}:=M\mathbf{v}$. Moreover, $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$ for $M\in\operatorname{EC}(d)$.
@@
The scalar matrices $zI$ with $\left|z\right|=1$ act trivially on projections, with $(zI\mathbf{v})(zI\mathbf{v})^{\dagger}=\mathbf{v}\mathbf{v}^{\dagger}$. The matrices $X$, $Z$, and $F$ all have entries in $\mathbb{Q}(\xi)$ (where $\xi=\xi_{d}=-e^{\frac{\pi i}{d}}$). The matrix $\sqrt{d}F$ also has entries in $\mathbb{Q}(\xi_{d})$, and $(F\mathbf{v})(F\mathbf{v})^{\dagger}=\frac{1}{d}(\sqrt{d}F\mathbf{v})(\sqrt{d}F\mathbf{v})^{\dagger}$. Therefore, writing any $M\in EC(d)$ as a product of the generators given in 3.7 shows that $(M\mathbf{v})(M\mathbf{v})^{\dagger}=M\mathbf{v}\mathbf{v}^{\dagger}M^{\dagger}$ has entries in $F^{\rm eproj}\!\left(S\right)$ (the compositum of $F^{\rm proj}\!\left(S\right)$ and $\mathbb{Q}(\xi)$). This argument shows that $F^{\rm proj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$, so $F^{\rm eproj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$. By considering $M^{-1}$, we also have $F^{\rm eproj}\!\left(S\right)=F^{\rm eproj}\!\left(M^{-1}MS\right)\subseteq F^{\rm eproj}\!\left(MS\right)$, so in fact $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right)$.
To show that $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$, it suffices to show that $\xi=-e^{-\frac{\pi i}{d}}$ occurs in the compositum of two fields $F^{\rm proj}\!\left(S\right)$ and $F^{\rm proj}\!\left(RS\right)$ where $R$ is the element of $\operatorname{C}(d)$ given in 3.7. Given a fiducial ${\mathbf{v}}$ generating $S=S([{\mathbf{v}}])$ having $v_{0}v_{1}\neq 0$, we consider the fiducial vector ${\mathbf{w}}=R{\mathbf{v}}$ of a geometrically equivalent SIC. It has entries $(R{\mathbf{v}})_{j}=\xi^{j(j+d)}v_{j}$ so $R{\mathbf{v}}_{0}=v_{0}$ and $R{\mathbf{v}}_{1}=-e^{\frac{\pi i}{d}}v_{1}$. Therefore $\Pi^{\prime}={\mathbf{w}}{\mathbf{w}}^{\dagger}$ has
$(\Pi^{\prime})_{1,0}=-e^{\frac{\pi i}{d}}v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(RS\right),$ (3.13)
while $v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. Now $v_{0}v_{1}\neq 0$ implies we may take their quotient in the compositum field and so get $-e^{\frac{\pi i}{d}}\in F^{\rm proj}\!\left(S\right)F^{\rm proj}\!\left(RS\right)$, as required. ∎
@@ 4. Algebraic SICs and the Galois action
We now focus on the subclass of algebraic SICs, for which the fields defined in Section 3.1 are algebraic number fields. In this paper, $\overline{\mathbb{Q}}$ is defined to be the algebraic closure of $\mathbb{Q}$ in $\mathbb{C}$, so there is a canonical choice of complex conjugation automorphism $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma_{c}(z)=\overline{z}$. We show that there is a well-defined Galois action on algebraic SICs and that the property of $S$ being an algebraic SIC is preserved by the subgroup of all automorphisms $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ that commute with $\sigma_{c}$ when restricted to the Galois closure of $F^{\rm proj}\!\left(S\right)$.
@@ 4.1. Algebraic SICs
An algebraic vector is some ${\mathbf{v}}\in{\overline{\mathbb{Q}}}^{d}$. For a projective-algebraic vector $\mathbf{v}$, we set its $\overline{\mathbb{Q}}$-equivalence class
$[[{{\mathbf{v}}}]]_{\operatorname{alg}}:=\{\lambda{\mathbf{v}}:\lambda\in\mathbb{C}\mbox{ and }\lambda{\mathbf{v}}\in\overline{\mathbb{Q}}^{d}\}.$ (4.1)
@@
Let ${\mathbf{v}}$ have some $v_{i}\neq 0$ and set $\mu=\frac{1}{v_{i}}$. Then ${\mathbf{w}}:=\mu{\mathbf{v}}=\left(\frac{v_{0}}{v_{i}},\frac{v_{1}}{v_{i}},\ldots,\frac{v_{d-1}}{v_{i}}\right)\in\overline{\mathbb{Q}}^{d}$ by the projective-algebraicity of ${\mathbf{v}}$. Now the complex conjugates $\sigma_{c}\!\left(\frac{v_{k}}{v_{i}}\right)=\frac{\overline{v}_{k}}{\overline{v}_{i}}$ are algebraic numbers, hence so is $\left|\frac{v_{k}}{v_{i}}\right|^{2}$. Thus, $\left|{\mathbf{w}}\right|=\sqrt{\sum_{k}\left|\frac{v_{k}}{v_{i}}\right|^{2}}$ is a positive real algebraic number. We set $\lambda=\frac{\mu}{\left|{\mathbf{w}}\right|}$ and obtain the required ${\mathbf{x}}=\frac{{\mathbf{w}}}{\left|{\mathbf{w}}\right|}=\lambda{\mathbf{v}}$ having $\left|{\mathbf{x}}\right|=1$.
Finally, if $\lambda$ and $\lambda^{\prime}$ give two such ${\mathbf{x}},{\mathbf{x}}^{\prime}\!$, necessarily $\lambda^{\prime}/\lambda\in\overline{\mathbb{Q}}$. ∎
@@
A $d$-dimensional line-SIC $S=\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ is algebraic if each of its $d^{2}$ vectors ${\mathbf{v}}_{i}$ are projective-algebraic.
@@
[…]
@@
    For an algebraic SIC the fields $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ are algebraic number fields; that is, finite algebraic extensions of $\mathbb{Q}$.
@@
Proof of (1): This follows using 4.2.
Proof of (2): That $F^{\rm vec}\!\left(S\right)$ is an algebraic number field follows directly from the definition of algebraic. Given a line-SIC specified by $d^{2}$ projective algebraic lines $\mathbb{C}{{\mathbf{v}}_{i}}$, using (1) we immediately have $F^{\rm trip}\!\left(S\right)$ being an algebraic number field. Furthermore, one may rescale all ${\mathbf{v}}_{i}$ to be algebraic unit vectors by 4.2. Now the associated projection matrices $\Pi_{{\mathbf{v}}}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$ have all entries in $\overline{\mathbb{Q}}$, hence $F^{\rm proj}\!\left(S\right)$ will be an algebraic number field.
Proof of (3): If the fiducial vector ${\mathbf{v}}$ of a Weyl–Heisenberg line-SIC $S$ is projective-algebraic, then by 4.2, ${\mathbf{v}}$ may be rescaled to be an algebraic unit vector ${\mathbf{w}}=\lambda{\mathbf{v}}$. The Weyl–Heisenberg group consists of matrices with algebraic entries, hence its action on ${\mathbf{w}}$ shows that all ${\mathbf{w}}_{i}$ for $1\leq i\leq d^{2}$ are algebraic, so $S$ is an algebraic line-SIC. The converse direction is immediate. ∎
It is conjectured that for each dimension $d\neq 3$, all Weyl–Heisenberg SICs are algebraic.
Conjecture 4.5 (Weyl–Heisenberg Algebraicity Conjecture).
For each dimension $d\geq 1$, except $d=3$, all Weyl–Heisenberg SICs are projective-algebraic.
This conjecture is equivalent to 2.5, which states that for all $d\neq 3$ the real projective algebraic variety $\operatorname{WHSIC}(d)$ is a finite set. It would follow by 4.4 that, if $S$ a Weyl–Heisenberg SIC, then $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$, and $F^{\rm proj}\!\left(S\right)$ are algebraic number fields.

### v3

@@ 3.2. Fields generated by Weyl–Heisenberg SICs
Let ${\mathbf{v}}=(v_{0},v_{1},\ldots,v_{d-1})^{\top}$ be a fiducial vector for $S$, and let $\Pi={\mathbf{v}}{\mathbf{v}}^{\dagger}$. For $X,Z$ as in 2.1, let $D_{{\mathbf{p}}}=X^{p_{1}}Z^{p_{2}}$ Then, for $0\leq p_{1}<d$ and $0\leq p_{2}<d$, the vectors ${\mathbf{v}}_{{\mathbf{p}}}=X^{p_{1}}Z^{p_{2}}{\mathbf{v}}$ form a set of generating set of vectors of the SIC, all having $\left|\left|\hskip 1.7pt{\mathbf{v}}_{{\mathbf{p}}}\hskip 1.7pt\right|\right|^{2}=\left|\left|\hskip 1.7pt{\mathbf{v}}\hskip 1.7pt\right|\right|^{2}$. Here ${\mathbf{p}}=(p_{1},p_{2})\in(\mathbb{Z}/p\mathbb{Z})^{2}$, and we also write ${\mathbf{v}}_{i}$ with $i=(d-1)p_{1}+p_{2}$ having $1\leq i\leq d^{2}$. We note for column vectors ${\mathbf{w}}_{i}$ that
$\langle{\mathbf{w}}_{1},{\mathbf{w}}_{2}\rangle\langle{\mathbf{w}}_{2},{\mathbf{v}}_{3}\rangle\langle{\mathbf{w}}_{3},{\mathbf{w}}_{1}\rangle$ $=\operatorname{Tr}\left({\mathbf{w}}_{1}^{\dagger}{\mathbf{w}}_{2}{\mathbf{w}}_{2}^{\dagger}{\mathbf{w}}_{3}{\mathbf{w}}_{3}^{\dagger}{\mathbf{w}}_{1}\right)$ (3.10)
$=\operatorname{Tr}\left({\mathbf{w}}_{1}{\mathbf{w}}_{1}^{\dagger}{\mathbf{w}}_{2}{\mathbf{w}}_{2}^{\dagger}{\mathbf{w}}_{3}{\mathbf{w}}_{3}^{\dagger}{\mathbf{w}}_{3}\right)$ (3.11)
$=\operatorname{Tr}\left(\Pi_{{\mathbf{w}}_{1}}\Pi_{{\mathbf{w}}_{2}}\Pi_{{\mathbf{w}}_{3}}\right)$ (3.12)
is a triple product, and similarly for $r$-fold cyclic scalar products of ${\mathbf{w}}_{i}$.
Proof of (1). The ratio of triple inner products
$\frac{\langle{\mathbf{v}},X^{-1}{\mathbf{v}}\rangle\langle X^{-1}{\mathbf{v}},Z{\mathbf{v}}\rangle\langle Z{\mathbf{v}},{\mathbf{v}}\rangle}{\langle{\mathbf{v}},Z^{-1}{\mathbf{v}}\rangle\langle Z^{-1}{\mathbf{v}},X{\mathbf{v}}\rangle\langle X{\mathbf{v}},{\mathbf{v}}\rangle}$ $=\frac{\langle X{\mathbf{v}},{\mathbf{v}}\rangle\langle{\mathbf{v}},XZ{\mathbf{v}}\rangle\langle Z{\mathbf{v}},{\mathbf{v}}\rangle}{\langle Z{\mathbf{v}},{\mathbf{v}}\rangle\langle{\mathbf{v}},ZX{\mathbf{v}}\rangle\langle X{\mathbf{v}},{\mathbf{v}}\rangle}$ (3.13)
$=\frac{\langle{\mathbf{v}},XZ{\mathbf{v}}\rangle}{\langle{\mathbf{v}},ZX{\mathbf{v}}\rangle}=\zeta_{d}.$ (3.14)
Via the SIC relations $\left|\langle{\mathbf{v}}_{i},{\mathbf{v}}_{j}\rangle\right|=\frac{1}{d+1}\left|{\mathbf{v}}\right|^{2}$, if $i\neq j$, and $\left|\left|\hskip 1.7pt{\mathbf{v}}\hskip 1.7pt\right|\right|^{2}$ if $i=j$, all factors in these ratios are nonzero. Therefore $\zeta_{d}\in F^{\rm trip}\!\left(S\right)$.
Proof of (2). Set $\nu_{\mathbf{p}}=\langle{\mathbf{v}},D_{\mathbf{p}}{\mathbf{v}}\rangle$. Set $\alpha_{1}:=\nu_{(1,0)}=\langle{\mathbf{v}},X{\mathbf{v}}\rangle$ and $\alpha_{2}=\nu_{(0,1)}=\langle{\mathbf{v}},Z{\mathbf{v}}\rangle$. Define the field $L(S):=F^{\rm trip}\!\left(S\right)(\alpha_{1},\alpha_{2})$. By (1), $\zeta_{d}\in F^{\rm trip}\!\left(S\right)$.
To show $F^{\rm proj}\!\left(S\right)\subseteq L(S)$, it suffices =to show that all entries $(\Pi_{i})_{k,\ell}\in L(S)$ for $1\leq i\leq d^{2}$ and $0\leq k,\ell<d$. Set ${\bf e}_{1}=(1,0)$ and ${\bf e}_{2}=(0,1)$. Now for ${\mathbf{p}}=p_{1}{\bf e}_{1}+p_{2}{\bf e_{2}}$, the product $\nu_{{\bf e}_{1}}^{p_{1}}\nu_{{\bf e}_{2}}^{p_{2}}\nu_{-{\mathbf{p}}}$ is a $(p_{1}+p_{2}+1)$-fold product (up to multiplication by a $d$-th root of unity), and thus by 3.3(2), $\nu_{{\bf e}_{1}}^{p_{1}}\nu_{{\bf e}_{2}}^{p_{2}}\nu_{-{\mathbf{p}}}\in F^{\rm trip}\!\left(S\right)(\alpha_{1},\alpha_{2},\zeta_{d}).$ We conclude that all $\nu_{-{\mathbf{p}}}$ are in $F^{\rm trip}\!\left(S\right)(\alpha_{1},\alpha_{2},\zeta_{d})$. The $D_{\mathbf{p}}$ form a basis of the matrix algebra ${\rm Mat}_{d\times d}(\mathbb{Q}(e^{\frac{2\pi i}{d}}))$. In particular all the matrices $E_{k,\ell}$ for $10\leq k,\ell\leq d-$ 1 with entry $1$ in the $(k,\ell)$ position and 00 elsewhere are linear combinations of the $D_{p}$ over the field $\mathbb{Q}(e^{\frac{2\pi i}{d}})$. Thus $(\Pi_{1})_{k,\ell}=\langle{\mathbf{v}},E_{k,\ell}{\mathbf{v}}\rangle\in F^{\rm trip}\!\left(S\right)(\alpha_{1},\alpha_{2})=L(S)$. Now for $i=(d-1)p_{1}+p_{2}$ we have
$(\Pi_{i})_{k,\ell}=\langle X^{p_{1}}Z^{p_{2}}{\mathbf{v}},E_{k,\ell}X^{p_{1}}Z^{p_{2}}{\mathbf{v}}\rangle=(\Pi_{1})_{k,\ell}$ (3.15)
holding for all $1\leq i\leq p^{2}$, by unitarity. Hence $(\Pi_{i})_{k,\ell}\in L(S)$ for all $i,k,\ell$. We conclude that $F^{\rm proj}\!\left(S\right)\subseteq L(S)$.
Finally, the product $\beta_{1}=(\alpha_{1})^{d}=\nu_{{\bf e}_{1}}^{d}=\nu_{{\bf e}_{1}}^{d}\nu_{0}$ is also a $(d+1)$-fold cyclic product and thus is in $F^{\rm inv}\!\left(S\right)=F^{\rm trip}\!\left(S\right)$, and similarly for $\beta_{2}=(\alpha_{2})^{d}=\nu_{{\bf e}_{2}}^{d}\in F^{\rm trip}\!\left(S\right)$. Thus $\alpha_{1}=\sqrt[d]{\beta_{1}}$ and $\alpha_{2}=\sqrt[d]{\beta_{2}}$ are both $d$-th roots of elements of $F^{\rm trip}\!\left(S\right)$. Since $\zeta_{d}\in F^{\rm trip}\!\left(S\right)$ we conclude adjoining these two elements to $F^{\rm trip}\!\left(S\right)$ gives an abelian extension $L(S)$ of $F^{\rm trip}\!\left(S\right)$, which is a Kummer extension whose degree divides $d^{2}$. The field $L(S)$ contains $F^{\rm proj}\!\left(S\right)$, and $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$, which implies $[F^{\rm proj}\!\left(S\right):F^{\rm trip}\!\left(S\right)]$ divides $d^{2}$.
@@
Proof of (3). By (1), since $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$, we have $\zeta_{d}\in F^{\rm proj}\!\left(S\right)$. Now $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)(e^{\frac{\pi i}{d}})$ is at most a quadratic extension of $F^{\rm proj}\!\left(S\right)$, since the field index $[\mathbb{Q}(e^{\frac{\pi i}{d}}):\mathbb{Q}(e^{\frac{2\pi i}{d}})]=2$ if $d$ is even and is $1$ if $d$ is odd. ∎
@@
where $X,Z$ are Weyl–Heisenberg group generators, $F$ is the discrete Fourier transform
$F_{j,k}=\frac{1}{\sqrt{d}}e^{\frac{2\pi ijk}{d}}\,\mbox{ for }\,0\leq j,k\leq d-1,$ (3.17)
and $R$ is a diagonal matrix with diagonal entries
@@
This is [46, Thm. 14.1]. Every group element is a product of finitely many generators, with exactly one use of a unitary scalar matrix $zI$. ∎
@@
Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$.
  • (1)
    For any $M\in\operatorname{EC}(d)$,
@@
    where $MS$ denotes the Weyl–Heisenberg line-SIC having fiducial vector ${\mathbf{w}}:=M\mathbf{v}$.
  • (2)
    $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$ for $M\in\operatorname{EC}(d)$.
@@
Proof of (1). To show $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right)$, note that scalar matrices $zI$ with $\left|z\right|=1$ act trivially on projections, with $(zI\mathbf{v})(zI\mathbf{v})^{\dagger}=\mathbf{v}\mathbf{v}^{\dagger}$. The matrices $X$, $Z$, and $F$ all have entries in $\mathbb{Q}(\xi)$ (where $\xi=\xi_{d}=-e^{\frac{\pi i}{d}}$). The matrix $\sqrt{d}F$ also has entries in $\mathbb{Q}(\xi)$, and $(F\mathbf{v})(F\mathbf{v})^{\dagger}=\frac{1}{d}(\sqrt{d}F\mathbf{v})(\sqrt{d}F\mathbf{v})^{\dagger}$. Writing any $M\in EC(d)$ as a finite product of the generators given in 3.7 shows that $(M\mathbf{v})(M\mathbf{v})^{\dagger}=M\mathbf{v}\mathbf{v}^{\dagger}M^{\dagger}$ has entries in $F^{\rm eproj}\!\left(S\right)$ (the compositum of $F^{\rm proj}\!\left(S\right)$ and $\mathbb{Q}(\xi)$). This argument shows $F^{\rm proj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$, so $F^{\rm eproj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$. By considering $M^{-1}$, we also have $F^{\rm eproj}\!\left(S\right)=F^{\rm eproj}\!\left(M^{-1}MS\right)\subseteq F^{\rm eproj}\!\left(MS\right)$, so $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right)$.
Proof of (2). To show $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$, it suffices to show that $\xi=-e^{-\frac{\pi i}{d}}$ occurs in the compositum of two fields $F^{\rm proj}\!\left(S\right)$ and $F^{\rm proj}\!\left(RS\right)$ where $R$ is the element of $\operatorname{C}(d)$ given in 3.7. Given a fiducial ${\mathbf{v}}$ generating $S=S([{\mathbf{v}}])$ having $v_{0}v_{1}\neq 0$, we consider the fiducial vector ${\mathbf{w}}=R{\mathbf{v}}$ of a geometrically equivalent SIC. It has entries $(R{\mathbf{v}})_{j}=\xi^{j(j+d)}v_{j}$ so $R{\mathbf{v}}_{0}=v_{0}$ and $R{\mathbf{v}}_{1}=\xi v_{1}$. Therefore $\Pi^{\prime}={\mathbf{w}}{\mathbf{w}}^{\dagger}$ has
$(\Pi^{\prime})_{1,0}=\xi v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(RS\right),$ (3.20)
while $v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. Now $v_{0}v_{1}\neq 0$ implies we may take their quotient in the compositum field and so get $\xi_{d}\in F^{\rm proj}\!\left(S\right)F^{\rm proj}\!\left(RS\right)$, as required. ∎
@@ 4. Algebraic SICs and the Galois action
We now treat the subclass of algebraic SICs, for which the fields defined in Section 3.1 are algebraic number fields. In this paper, $\overline{\mathbb{Q}}$ is defined to be the algebraic closure of $\mathbb{Q}$ inside $\mathbb{C}$, so there is a canonical choice of complex conjugation automorphism $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma_{c}(z)=\overline{z}$. We show that there is a well-defined Galois action on algebraic SICs and that the property of $S$ being an algebraic SIC is preserved by the subgroup of all automorphisms $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ that commute with $\sigma_{c}$ when restricted to the Galois closure of $F^{\rm proj}\!\left(S\right)$.
@@ 4.1. Algebraic SICs
An algebraic vector is any ${\mathbf{v}}\in{\overline{\mathbb{Q}}}^{d}$. For a projective-algebraic vector $\mathbf{v}$, its $\overline{\mathbb{Q}}$-equivalence class is
$[[{{\mathbf{v}}}]]_{\operatorname{alg}}:=\left\{\lambda{\mathbf{v}}:\lambda\in\mathbb{C}\mbox{ and }\lambda{\mathbf{v}}\in\overline{\mathbb{Q}}^{d}\right\}.$ (4.1)
@@
Let ${\mathbf{v}}$ have some $v_{i}\neq 0$ and set $\mu=\frac{1}{v_{i}}$. Then ${\mathbf{w}}:=\mu{\mathbf{v}}=\left(\frac{v_{0}}{v_{i}},\frac{v_{1}}{v_{i}},\ldots,\frac{v_{d-1}}{v_{i}}\right)\in\overline{\mathbb{Q}}^{d}$ by the projective-algebraicity of ${\mathbf{v}}$. Now the complex conjugates $\sigma_{c}\!\left(\frac{v_{k}}{v_{i}}\right)=\frac{\overline{v}_{k}}{\overline{v}_{i}}$ are algebraic numbers, hence so is $\left|\frac{v_{k}}{v_{i}}\right|^{2}$. Thus, $\left|{\mathbf{w}}\right|=\sqrt{\sum_{k}\left|\frac{v_{k}}{v_{i}}\right|^{2}}$ is a positive real algebraic number. We set $\lambda=\frac{\mu}{\left|{\mathbf{w}}\right|}$ and obtain the required ${\mathbf{x}}:=\lambda{\mathbf{v}}=\frac{{\mathbf{w}}}{\left|{\mathbf{w}}\right|}$, having $\left|{\mathbf{x}}\right|=1$.
Finally, if $\lambda$ and $\lambda^{\prime}$ give two algebraic ${\mathbf{x}},{\mathbf{x}}^{\prime}\!$, necessarily $\lambda^{\prime}/\lambda\in\overline{\mathbb{Q}}$. ∎
@@
A $d$-dimensional line-SIC $S=\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ is algebraic if each of its $d^{2}$ vectors ${\mathbf{v}}_{i}$ are projective-algebraic (or equivalently, if $F^{\rm vec}\!\left(S\right)$ is a finite extension of $\mathbb{Q}$).
@@
    A line-SIC is algebraic if and only if, for unit vectors ${\mathbf{v}}_{i}$, all the SIC projection matrices $\{\Pi_{{\mathbf{v}}_{i}}:1\leq i\leq d^{2}\}$ have algebraic number entries.
@@
    For an algebraic SIC $S$, the fields $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$, $F^{\rm proj}\!\left(S\right)$ and $F^{\rm eproj}\!\left(S\right)$ are algebraic number fields, that is, finite algebraic extensions of $\mathbb{Q}$.
@@
Proof of (1). This follows using 4.2.
Proof of (2). That $F^{\rm vec}\!\left(S\right)$ is an algebraic number field follows directly from the definition of algebraic. Given a line-SIC specified by $d^{2}$ projective algebraic lines $\mathbb{C}{{\mathbf{v}}_{i}}$, using (1) we immediately have $F^{\rm trip}\!\left(S\right)$ being an algebraic number field. Furthermore, one may rescale all ${\mathbf{v}}_{i}$ to be algebraic unit vectors by 4.2. Now the associated projection matrices $\Pi_{{\mathbf{v}}}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$ have all entries in $\overline{\mathbb{Q}}$, so $F^{\rm proj}\!\left(S\right)$ is an algebraic number field. Finally, $F^{\rm eproj}\!\left(S\right)$ is a finite extension of $F^{\rm proj}\!\left(S\right)$.
Proof of (3). If the fiducial vector ${\mathbf{v}}$ of a Weyl–Heisenberg line-SIC $S$ is projective-algebraic, then by 4.2, ${\mathbf{v}}$ may be rescaled to be an algebraic unit vector ${\mathbf{w}}=\lambda{\mathbf{v}}$. The Weyl–Heisenberg group $\operatorname{WH}(d)$ consists of matrices with algebraic entries, hence its action on ${\mathbf{w}}$ shows that all ${\mathbf{w}}_{i}$ for $1\leq i\leq d^{2}$ are algebraic, so $S$ is an algebraic line-SIC. The converse direction is immediate. ∎